### v1

# Andreev spin qubit protected by Franck-Condon blockade

Abstract
Andreev levels localized in a weak link between two superconductors can trap a superconducting quasiparticle. If there is a spin-orbit coupling in the link, the spin of the quasiparticle couples to the Josephson current. This effect can be leveraged to control and readout the spin of the quasiparticle thus using it as a qubit. One of the factors limiting the performance of such an Andreev spin qubit is spin relaxation. Here, we theoretically demonstrate that the relaxation lifetime can be enhanced by utilizing the coupling between the Andreev spin and the supercurrent in a transmon circuit. The coupling ensures that the flip of the quasiparticle spin can only happen if it is accompanied by the excitation of multiple plasmons, as dictated by the Frank-Condon principle. This blocks spin relaxation at temperatures small compared to plasmon energy.
Introduction. The spin of a single quasiparticle trapped in a Josephson junction can couple to the supercurrent via the spin-orbit interaction [1]. Following several years of theoretical work [1, 2, 3, 4, 5, 6, 7], this effect has been experimentally established via microwave spectroscopy measurements [8, 9]. It is a key element for the operation of Andreev spin qubits [10, 11, 12]. These novel qubits have multiple properties favorable for quantum computing: a $\mu\mathrm{m}$-scale spatial footprint, large relative anharmonicity, and a direct interaction with microwave fields mediated by the spin-supercurrent coupling. The latter coupling also allows for cm-scale qubit connectivity [12] which is useful for quantum simulations [13].
In the current experimental realizations of Andreev spin qubits, however, the benefits listed above are outweighed by short decoherence times [11, 14, 15]. Both dephasing and relaxation times of Andreev spin qubits are much shorter than those of state-of-the-art semiconductor spin qubits [16, 17]. At least in part, these times could be improved by updating the material platform used in Andreev spin qubit devices. So far, all measured devices were based on InAs/Al nanowires, which are far from optimal for spin coherence [16]. Material considerations aside, one may wonder: in the spirit of hardware-protected qubits [18], is it possible to enhance the coherence of the Andreev spin qubit by circuit design?
Figure 1: Protected Andreev spin qubit. Left: A spin-$1/2$ quasiparticle (green) trapped inside a Josephson junction forming a resonant transmon circuit. The quasiparticle spin couples to the supercurrent flowing in the circuit via the spin-orbit coupling, leading to a spin-dependent Josephson potential $U_{\pm}(\phi)$. The noise in the magnetic environment of the spin (dark yellow spins) causes it to relax. The quasiparticle spin can also be externally driven (blue arrow), e.g. via a local gate. Right: The potential energy of the circuit splits in two branches (dashed lines), one for spin up (blue) and one for spin down (red). The transmon ground state wave functions for opposite spins are disjoint, which protects the spin from relaxation. The protection can be compromised by a finite temperature, which activates transitions where circuit plasmons are excited in addition to the spin flip (vertical arrow). By the Franck-Condon principle, the matrix elements for such transitions are larger than those without plasmons (horizontal arrow).
In this Letter, we show that the relaxation time of the Andreev spin qubit can be enhanced by shunting it with a capacitor, as in a transmon circuit [19]. Then, provided the spin-dependent supercurrent is large enough, the transmon ground state wave-functions for opposite spin orientations become disjoint, separated by a distance $\sim 2\phi_{0}$ larger than their extent $\sim 2\phi_{c}$, see Fig. 1. In this situation, the matrix elements for spin-flip transitions are exponentially suppressed with $(\phi_{0}/\phi_{c})^{2}$, leading to an enhancement of the spin relaxation time. The suppression of transition matrix elements is governed by Franck-Condon principle [20], and we therefore refer to it as the Franck-Condon blockade, by analogy with a related phenomenon in molecular transport [21, 22]. The pure dephasing time of the Andreev spin qubit is neither improved nor worsened by Frank-Condon blockade.
We build a systematic theory of Frank-Condon blockade of Andreev spin qubit relaxation. We describe how the protection can be compromised by finite temperature. In fact, we show that the residual spin relaxation comes from a thermally-activated process where a spin-flip is accompanied by an excitation of several plasmonic quanta. Despite requiring energy, this process has a larger matrix element compared to the excitation-less process, as described by the Frank-Condon principle. Additionally, we identify experimentally observable consequences of the Frank-Condon blockaded regime: the appearance of plasmonic transitions in response to an external drive acting on the spin and a staircase-shaped dependence of spin lifetime on magnetic field. Finally, we provide a guideline for future experiments by discussing the device design of the protected Andreev spin qubit. Throughout the paper we use units with $k_{B}=\hbar=1$.
Model and main idea. We consider a circuit in which the junction trapping the spin-$1/2$ quasiparticle is shunted by capacitance $C$, see Fig. 1. Adopting a minimal model for the Josephson energy [5], the Hamiltonian of the circuit is given by
$H=4E_{c}(i\partial_{\phi}+n_{g})^{2}+E_{0}\cos(\phi)+E_{\textrm{so}}\sigma_{z}\sin(\phi)\,.$ (1)
Here, $\phi$ is the superconducting phase difference, while $\sigma_{x,y,z}$ are spin Pauli matrices. $E_{c}=e^{2}/2C$ is the single-electron charging energy, and $n_{g}$ is a shift due to stray offset charges. The couplings $E_{0}$ and $E_{\textrm{so}}$ represent the spin-independent and spin-dependent contributions to the amplitude of Cooper pair tunneling across the junction in the presence of the trapped quasiparticle. This model has been applied successfully to quantum dot Josephson junctions in InAs/Al devices [23].
The last term of Eq. (1) arises due to spin-orbit interaction. It is the simplest coupling between spin and superconducting phase compatible with time-reversal symmetry (which sends $\phi\to-\phi$ and $\sigma_{z}\to-\sigma_{z}$) and the compactness of $\phi$. In principle, the direction $\hat{n}$ of the spin-orbit field may depend on the phase difference, so that $\sigma_{z}$ should be replaced by a more general operator $\vec{\sigma}\cdot\hat{n}(\phi)$ with $\hat{n}(\phi)=\hat{n}(-\phi)$. This replacement opens new channels of spin relaxation, as it couples the spin to charge fluctuations via $n_{g}$. This, however, does not qualitatively affect our results regarding the Frank-Condon blockade. We therefore resort to the simple form of the coupling Hamiltonian given by Eq. (1).
The potential in Eq. (1) consists of two sinusoidal Josephson energy branches of equal amplitude $\sqrt{E_{0}^{2}+E_{\textrm{so}}^{2}}$, see Fig. 1. The energy minima occur at $\phi=\pi\pm\phi_{0}$ for spin up and down states respectively, with
$\phi_{0}=\arctan(E_{\textrm{so}}/E_{0})\,.$ (2)
The size of the quantum fluctuations is set by
$\phi_{c}=\sqrt{\frac{4E_{c}}{\omega_{p}}}\,,$ (3)
where $\omega_{p}=\sqrt{8E_{c}(E_{0}^{2}+E_{\textrm{so}}^{2})^{1/2}}$ is the Josephson plasma frequency of the small oscillations of the phase around the minima. Each energy level in the system is Kramers-degenerate due to the presence of the spin degree of freedom. When $E_{c}\ll\omega_{p}$, the Josephson potential can be approximated by a parabola and the low-lying levels are equidistant, $E_{k}=\omega_{p}(k+\tfrac{1}{2})$, where the integer number $k$ is the number of plasmon excitations. In this limit, one can neglect the exponentially small dependence of energy levels on $n_{g}$. In a more accurate treatment, which is beyond the scope of the manuscript, the $E_{k}$ are the energy levels of the Cooper-pair box Hamiltonian [19].
To describe Frank-Condon blockade, we introduce a parameter $\xi_{0}=(\phi_{0}/\phi_{c})^{2}$. If $\xi_{0}~{}\gg~{}1$, the lowest-level wave functions with opposite spin are disjoint. Their overlap is exponentially small in $\xi_{0}$, since it is the integral of a product of two Gaussian wave functions of width $\sim\phi_{c}$ at a distance $\sim\phi_{0}$, see Fig. 1. We call regime $\xi_{0}\gg 1$ “the strong coupling regime”, referring to the strength of coupling between the spin and the supercurrent. To achieve the strong coupling regime it is sufficient to fulfill two conditions: (i) $\omega_{p}\gg E_{c}$ such that the circuit is in the transmon regime [19] with small zero-point phase fluctuations ($\phi_{c}\ll 1$), (ii) $E_{\textrm{so}}\gtrsim E_{0}$ such that the separation between potential minima for opposite spin is large ($\phi_{0}\sim 1$).
Figure 2: (a): Temperature dependence of the spin-flip rate at $J_{z}=0$. We assume that the system is initialized in the plasmon ground state $k=0$ for one of the two spins. The solid line is Eq. (6), the purple dashed line is the approximate expression (8), and the horizontal dashed line is the $T=0$ result (5). The parameters are $E_{c}/E_{\textrm{so}}=0.035$ and $E_{0}/E_{\textrm{so}}=1$, yielding $\omega_{p}/E_{\textrm{so}}\approx 0.63$ and $\xi_{0}\approx 2.8$. (b)-(c): Dependence of the spin-flip rate on temperature and magnetic field parallel to the spin-orbit axis, starting from either spin down (b) or spin up (c). The rates display step-like features as a function of the magnetic field, which are washed out by increasing temperature. The vertical and horizontal lines correspond to the line cuts shown in panels (a) and (d) respectively. (d): Magnetic field dependence of the spin-flip rate for initial states with opposite spin (initialized in a state with $k=0$). The solid curves are computed at $T/\omega_{p}=0.1$, while the dashed ones at $T/\omega_{p}=1$. In the low temperature curves, the steps occur at half-integer values of $J_{z}/\omega_{p}$, and have heights given by the coefficients $w_{k}$. The results presented in this figure are derived for a simple bath of two-level systems, with a constant density of states and energy-independent coupling to the Andreev spin.
Assume now that the Andreev spin is coupled to a vector field $\vec{J}$, i.e., we add an extra term $H_{J}=\vec{J}\cdot\vec{\sigma}$ to the Hamiltonian in Eq. (1). This vector field can describe either an intentionally applied magnetic field or the uncontrolled noise stemming from the environment of the spin. To demonstrate the essence of the Frank-Condon blockade, we can momentarily leave the origin of $\vec{J}$ unspecified. If we project $H_{J}$ onto the $k=0$ subspace, we obtain the effective Hamiltonian:
$H_{\textrm{eff}}=e^{-\xi_{0}/2}(J_{x}\sigma_{x}+J_{y}\sigma_{y})+J_{z}\sigma_{z}\,.$ (4)
We see that the matrix elements of $H_{J}$ between ground states of opposite spins are suppressed a factor $e^{-\xi_{0}/2}$ which is small when $\xi_{0}\gg 1$. Thus, the external field is not effective in flipping the Andreev spin, and the qubit encoded in this spin is protected from relaxation. This is the main idea of our proposal.
Before assessing the conditions for the validity of Eq. (4), let us show explicitly that it leads to the enhancement of spin relaxation. To this end, we assume that the field ${\vec{J}}$ encodes a spin-flipping environment of the Andreev spin. Realistically, such environment will be constituted by nuclear spins in combination with phonons [24], exchanged-coupled itinerant quasiparticles in the superconducting leads [25], magnetic field noise, and charge noise. However, on a qualitative level, the nature of the environment is not important for the discussion of Frank-Condon blockade. Therefore, to be concrete, we adopt a simplified model in which the environment is a collection of two-level systems of an unspecified origin. We assume that this bath has a constant density of states $\nu_{0}$ and each two-level system couples to $\sigma_{x}$ with the same strength $g$. Using Fermi’s golden rule, we then find
$\Gamma_{\uparrow\rightarrow\downarrow}=\Gamma_{0}e^{-\xi_{0}},\quad\Gamma_{0}=2\pi\nu_{0}|g|^{2}.$ (5)
As $\xi_{0}$ is increased, the spin relaxation time increases exponentially. We now explore limitations to this result.
Temperature dependence of spin relaxation rate. The effective Hamiltonian given by Eq. (4) does not take into account the plasmonic excitations (states with $k>0$). At the same time, the full version of the coupling Hamiltonian Hamiltonian $H_{J}$ mixes the low-energy states, $k=0$, with the excited states, $k>0$. For this reason, when describing the dynamics of the Andreev spin, neglecting the excited states is only allowed when the relevant energy scales (such as the temperature or $\vec{J}$) are much smaller than $\omega_{p}$. Otherwise, a more careful calculation of the spin-flip dynamics is warranted. In fact, as we now demonstrate, a temperature as low as $\omega_{p}/\ln(\xi_{0})$ can significantly compromise the protection.
To reveal the role of finite temperature, we resort to the two-level system model for the bath described above. For simplicity, we assume that the system is initialized in the plasmon ground state, $k=0$, for one of the spin directions. The total spin-flip rate can then be evaluated as
$\Gamma_{\uparrow\rightarrow\downarrow}=2\Gamma_{0}\sum_{k}\frac{w_{k}}{1+e^{\beta\omega_{p}k}}\,,\;\;w_{k}=\frac{e^{-\xi_{0}}\xi_{0}^{k}}{k!}\,,$ (6)
with $\beta=1/T$. The weights $w_{k}\equiv\left|g_{k0}\right|^{2}$ originate from the Franck-Condon overlap factor $g_{kk^{\prime}}$, defined as
$g_{kk^{\prime}}=\left\langle k\!\downarrow\right|\sigma_{x}\left|k^{\prime}\!\uparrow\right\rangle=\int_{-\infty}^{\infty}\psi_{k}(\phi+\phi_{0})\,\psi_{k^{\prime}}(\phi-\phi_{0})d\phi\,,$ (7)
where $\psi_{k}$ are the wave functions of the harmonic oscillator. At small temperature $T\ll\omega_{p}$, factors $(1+e^{\beta\omega_{p}k})^{-1}$ in Eq. (6) can be approximated as $e^{-\beta\omega_{p}k}\ll 1$ for any $k>0$. Then, we find
$\Gamma_{\uparrow\rightarrow\downarrow}=\Gamma_{0}e^{-\xi_{0}}\left(2e^{\xi_{0}e^{-\beta\omega_{p}}}-1\right).$ (8)
According to this equation, Eq. (5) is accurate as long as the temperature satisfies $T\ll\omega_{p}/\ln\xi_{0}$. Otherwise, Eq. (5) underestimates the transition rate. The rate is enhanced due to the possibility of a spin-flip with a simultaneous excitation of $k$ plasmons. Even though this process is penalized by a factor $e^{-\beta\omega_{p}k}$, it has a much larger matrix element compared to that unaccompanied by plasmon excitations. In the regime $T\gg\omega_{p}$, we obtain $\Gamma_{\uparrow\rightarrow\downarrow}=\Gamma_{0}$ since $e^{\beta\omega_{p}k}\approx 1$ for all $k$ and the following sum rule holds:
$\sum_{k}w_{k}=e^{-\xi_{0}}\sum_{k}\frac{\xi_{0}^{k}}{k!}=1\,.$ (9)
Finally, we note that Eq. (6) can be extended to the case in which the plasmons are thermally occupied (as opposed to taking $k=0$ as the initial state). This brings no qualitative change to the conclusions of this section.
Effect of a static magnetic field. Next, we discuss experimental signatures of a protected regime of the Andreev spin qubit for $T\ll\omega_{p}$. The first such signature is the unusual magnetic field dependence of the spin-flip rate. To see this dependence, we assume that a static magnetic field is applied parallel to the spin-orbit axis and so we include a term $J_{z}\sigma_{z}$ in the Hamiltonian. Then, Eq. (6) generalizes to
$\Gamma_{\uparrow\rightarrow\downarrow}=2\Gamma_{0}\sum_{k}\frac{w_{k}}{1+e^{\beta\omega_{p}k}e^{-2\beta J_{z}}}.$ (10)
[…]
The Franck-Condon factors of Eq. (7) also determine the splitting of Kramers-degenerate excited states due to magnetic field perpendicular to the spin-orbit axis. Indeed, according to first-order perturbation theory, term $H_{J}=J_{x}\sigma_{x}$ in the Hamiltonian splits the energy of the $k$-th level by $\delta E_{k}\approx J_{x}g_{kk}=J_{x}e^{-\xi_{0}/2}L_{k}(\xi_{0})$, with $L_{k}$ the $k$-th Laguerre polynomial.
[…]
$\Omega_{k0}=J_{x}\left|g_{k0}\right|=J_{x}e^{-\xi_{0}/2}\xi_{0}^{k/2}/\sqrt{k!}\,.$ (11)
The energy levels excited by the drive may be broadened, such that the Rabi oscillations are over-damped. In this case, the spectrum should be characterized with a rate of incoherent spin-flip transitions at different drive frequencies. According to Fermi’s golden rule, the rate is given by
$\Gamma_{\uparrow\to\downarrow}$ $=2\pi J_{x}^{2}\,\sum_{k=0}^{\infty}\frac{w_{k}}{\pi}\frac{\Gamma}{(\omega-k\omega_{p}-2J_{z})^{2}+\Gamma^{2}}\,,$ (12)
where $\Gamma$ is a parameter describing the broadening. The resulting rate as a function of the spin-orbit parameter $E_{\mathrm{so}}$ is shown in Fig. 3.
In the absence of spin-orbit coupling ($E_{\textrm{so}}=0$, and thus $\xi_{0}=0$), one has $w_{0}=1$ and $w_{k>0}=0$. Therefore, as is apparent from Fig. 3, the only active transition is the spin-flip transition with $k=0$. At finite $E_{\textrm{so}}$, instead, the absorption spectrum consists of a set of distinct transitions involving the excitation of $k$ plasmons. Indeed, the sum rule of Eq. (9) guarantees that the exponential suppression of $w_{0}$ is compensated by a power-law increase in the strength of higher-frequency transitions. If $\xi_{0}\gg 1$, using Stirling’s approximation, one can see that the maximum transition strength occurs for $k=\lfloor\xi_{0}\rfloor$. Therefore, in the strong coupling regime, it is easier to excite a spin-flip transition if it is accompanied by the excitation of multiple plasmons, because of a larger wave function overlap with such a final state.
Finally, we note that although we considered a purely magnetic drive, it is also possible to coherently drive the spin via a gate, using electron dipole spin resonance [27, 28, 29].
Figure 3: Spin-flip rate in response to an external drive at frequency $\omega$ as a function of the coupling $E_{\textrm{so}}$, using Eq. (12). We set $T=0$, $J_{z}/2\pi=0.2$ GHz, $E_{0}/2\pi=1$ GHz, $E_{c}/2\pi=35$ MHz, $\Gamma=50$ MHz. The results illustrate that in the presence of spin-orbit coupling, the spin-flip transition fans out into a set of plasmonic transitions with spacing $\omega_{p}$.
Experimental considerations. Demonstrating the Frack-Condon blockade requires the realization of a strong coupling regime, $\xi_{0}\gg 1$, which is within reach of current experiments. Indeed, measurements of Andreev spins in InAs nanowires routinely observe $E_{0}/2\pi\sim E_{\textrm{so}}/2\pi\sim 1~{}\mathrm{GHz}$ [8, 9, 23, 14]. In that case, to reach strong coupling it is sufficient to construct a circuit with a small charging energy, $E_{C}/2\pi\ll 1~{}\mathrm{GHz}$. Transmon circuits with values as low as $36~{}\mathrm{MHz}$ have been engineered [30]. For these values, $\xi_{0}\approx 3$ can be expected. Potentially, $\xi_{0}$ can be further increased by finding material platforms leading to higher values of $E_{\textrm{so}}$.
Another challenge is the requirement $T\ll\omega_{p}$, but we believe that this condition can also be met in the lab. For the parameters listed above, we expect $\omega_{p}/2\pi\sim 600~{}\mathrm{MHz}$, that is $\sim 30~{}\mathrm{mK}$, above the $\sim 10~{}\mathrm{mK}$ achievable by modern dilution refrigerators. The feasibility of operating low-frequency transmons with small thermal populations was experimentally demonstrated in Ref. [30].
Two final requirements for the operation of the protected Andreev spin qubit concern the feasibility of qubit readout and gates. The spin can be read out dispersively by coupling the circuit to a readout resonator [10, 23]. To achieve such a spin readout, time-reversal symmetry has to be broken during the measurement, for instance by applying a magnetic field perpendicular to spin-orbit field or by threading magnetic flux through a SQUID loop  [^1] See e.g. the Supplemental Information of Ref. [23] for a demontration of Andreev spin readout via a resonator.. The realization of single-qubit gates is complicated by the suppressed overlap between the transmon ground state wave functions for different spins. Fortunately, the Rabi frequency scales with $\xi_{0}$ more mildly, as $e^{-\xi_{0}/2}$, than relaxation, as $e^{-\xi_{0}}$. We leave a detailed analysis of single and two-qubit gates for future investigation.
Conclusions. We have shown that an Andreev spin qubit can be protected against spin relaxation by shunting it with a large capacitance. We provided a set of specific experimental predictions that distinguish the protected regime from an unprotected one. These include an unusual dependence of the spin relaxation rate on the magnetic field as well as the appearance of additional branches detectable in the two-tone spectrum of the system.
We have highlighted the limitations on the protected regime posed by finite temperature of the device. This thermal fragility can be related to the Franck-Condon principle via the sum rule of Eq. (9). The exponential suppression of spin flips due to the disjoint support of the low-lying plasmon states is accompanied by the simultaneous increase of the matrix elements involving the excited plasmon states. At finite temperature, this diminishes the expected returns of being in a protected regime. A similar trade-off could apply to any protected qubit based on a double-well energy landscape [18].
The protected Andreev spin qubit that we propose has noise strongly biased towards pure dephasing. Such a noise bias may be advantageous for quantum error correction [32, 33]. Along these lines, it will be interesting to merge the ideas presented in our manuscript with those of the recent blueprint for Kramers-protected error correction in Andreev spin qubit arrays [34]. Other extensions are the inclusion of a superinductor [35] into the circuit and the study of the Franck-Condon blockade in multi-terminal junctions [36, 37, 38].
Acknowledgments. We thank Valla Fatemi, Lukas Splitthoff, Spencer Diamond, Heekun Nho, Vladislav Kurilovich, Christian K. Andersen, Arno Bargerbos, Marta Pita-Vidal, Jaap Wesdorp, Maximilian Rimbach-Russ, Aleksandr Svetogorov, Leandro Tosi, Leonid Glazman, and Michel Devoret for useful discussions and feedback on the manuscript.
The research of P.D.K., T.C., and C.G.L.B. was sponsored by the Army Research Office (ARO) under grants no. W911NF-22-1- 0053 and W911NF-23-1-0051, by DARPA under grant no. HR0011-24-2-0346, by the U.S. Department of Energy (DoE), Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under contract number DE-SC0012704. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the ARO, DARPA, DoE, or the US Government. The US Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
Data availability statement. All data generated and code used in this work are available on Zenodo.